# Uranus and Neptune as methane planets: producing icy giants from refractory planetesimals

### v1

@@ 3.4 Models with Methane
We also consider a third ice, methane, as motivated by the following arguments. The CH abundance in the upper atmosphere of Uranus and Neptune is observed to be far greater than solar (Atreya et al., 2020). This should not be regarded as indicative of the bulk abundance of methane, since these measurements only probe the atmosphere at the 1 bar level. Nevertheless, early studies did in fact consider models where CH was mixed into the entire envelope with far greater abundances than solar (Podolak, 1976; Podolak and Reynolds, 1981), or with an icy layer that consisted primarily of HO and CH (Hubbard and Macfarlane, 1980). This was motivated by the idea that nebular carbon should condense to form CH and accrete onto Uranus and Neptune in large amounts. At around the same time, however, the view was shifting away from condensation of solid CH as a result of kinetic inhibition (Lewis and Prinn, 1980). Instead, at sufficiently low temperatures, the expectation was that the large majority of carbon would condense directly to CO (Lewis and Prinn, 1980). Later it was briefly suggested that, instead, CO might be chemically transformed in the atmosphere of the gas giants and convert to CH (Podolak and Reynolds, 1984) when reacting with hydrogen, altering their bulk composition. However, this hypothesis did not develop further since. Here we extend the idea that CH could be a key volatile in the outer planets. As already outlined in Section 2, while CO is almost negligible and therefore has a very small potential contribution for generating methane in typical contemporary planetesimals, refractory organics contribute to a very significant portion of their mass. These refractory organics may also chemically transform into copious amounts methane when they accrete in vast quantities onto the planet’s atmosphere, and are capable of altering its bulk composition.
We use the same setup as in the previous section, but now replace carbon monoxide with methane. When we incorporate CH as our second icy material, the results are very different. Randomly generated interior models of Uranus and Neptune with CH are shown in Figure 3. In these models the water/rock mass ratio now spans the range 0.1-10. As noted above, CO as an additional ice tends to replace SiO. CH, on the other hand, tends to replace water, so the rock/water mass ratio can be significantly larger.
@@
Finally, in Figure 5, we show the methane mass fraction as a function of the water/rock mass ratio. The methane mass fraction probed by the space of possible solutions is at least 10%. In models that also have a large rock/water mass fraction (left of the vertical dashed line) it is necessarily larger, upwards of 20%. Since the envelope mass fraction in these models also has large values (Figure 3), the combination of large envelope and even larger CH mass fractions means that the other materials - HO, SiO, and Fe, are collectively only a relatively small fraction of the planet’s mass. Thus, in models that have a large rock/water mass fraction, specifically, the methane fraction might even be larger than that of HO, SiO, and Fe combined. This is a difficulty because CH is certainly not that prevalent in the contemporary solar system, and is normally a very small fraction compared to water (Bockelée-Morvan and Biver, 2017). Thus, for methane to participate in solving the composition tension problem, a large methane fraction must be justified. We suggest that organic-rich refractories are sufficiently abundant in outer solar system planetesimals to drive chemical reactions in the atmospheres of Uranus and Neptune, which, during the phase of planet growth, may produce the required methane. These processes are discussed next.
@@ 4 Growth of the Outer Planets and Accretion of Solids
We consider a growing planet in the outer solar system, which starts accreting mass onto an embryo-sized core (the incipient core mass in the range 0.05-0.5 M (Brouwers et al., 2018; Bodenheimer et al., 2018)). Classical models emphasized subsequent growth due to accretion of large planetesimals. These planetesimals could fall directly onto the core (Podolak et al., 1988) however they could also dissolve in the atmosphere due to thermally (Iaroslavitz and Podolak, 2007; Brouwers et al., 2018) or mechanically (Mordasini et al., 2015) driven ablation. As core mass accumulates, additional atmosphere is collected (Pollack et al., 1996). Energy release is added to the luminosity of the planet and thus internal temperatures increase during growth (Pollack et al., 1996).
@@
We can get an idea of the relevant density profile, as well as the temperature-pressure (T-P) profiles, using a newly updated code (Stevenson et al., 2022) implemented here for the first time for a growing Uranus. This code solves the equations of mass conservation, hydrostatic equilibrium, energy conservation, and energy transport assuming spherical symmetry. It includes the effects of mass and energy deposition by accreting solids and gases, and equations of state of water, SiO and iron. It also includes opacity due to dust grains, the ablation and breakup of (here only 100 km mono-sized) solid planetesimals in the gaseous envelope, treatment of compositional gradients and heating by the host star which drives mass loss. Using this code, we obtain snapshots of the density/temperature/pressure profiles of the atmosphere, at three different times after the start of the evolution: 1.35, 2.06 and 2.65 Myr.
We then use these atmospheric density profiles as input to an ablation code (Podolak et al., 1988), which includes the detailed effects of frictional gas drag, heating by ambient thermal radiation, and the ensuing evaporation (sublimation) erosion derived from the energy input. Following the trajectories of accreted planetesimals of various sizes, we can determine the characteristic temperature-pressure conditions during ablation. For the parameters we selected, we find that small sub-km planetesimals are strongly slowed by atmospheric drag and ablate as a consequence. The larger planetesimals are correspondingly less affected by drag and are therefore able to maintain the speed necessary to break up deeper in the atmosphere, due to ram pressure. As already mentioned, these results depend on the assumed material properties. Since the code is unable to account for the intricacies of planetesimal intra-porosity or material mixing, we employ a similar approach to other studies (Iaroslavitz and Podolak, 2007; Brouwers et al., 2018; Brouwers and Ormel, 2020; Ormel et al., 2021) and simplify matters by considering non-porous planetesimals with a uniform density of 2000 kg m, and the material thermodynamic/cohesion properties of water, which might represent a mean value for a mixture of silicates, organics, water and other high-volatility ices. Given these assumptions, the temperature-pressure conditions during ablation are provided in Table 1. Note that ablation is a gradual process and therefore it occurs for a range of T-P conditions. To simplify matters, for sub-km planetesimal we listed a single T-P value that matches the moment during which half of the planetesimal mass has eroded. For larger planetesimals, the listed value matches the moment of ram-pressure breakup when most of the mass is instantantly delivered to the atmosphere.
@@ 4.1 Compositional Stratification During Accretion
Once high-Z materials are present in the atmosphere, they are expected to form a compositional gradient inside the planet (Helled and Stevenson, 2017; Lozovsky et al., 2017; Brouwers et al., 2018; Bodenheimer et al., 2018; Valletta and Helled, 2019; Vazan and Helled, 2020; Brouwers and Ormel, 2020; Ormel et al., 2021). Indeed there is some seismic and gravity-based evidence for a deep compositional gradient in Saturn (Fuller, 2014; Mankovich and Fuller, 2021) and Jupiter (Wahl et al., 2017; Debras and Chabrier, 2019).
Assuming that high-Z vapor in the atmosphere obeys the ideal gas law, an assumption which was found to be judicious in giants such as Uranus and Neptune (Iaroslavitz and Podolak, 2007), then compositional stratification can be explained as follows. At a given layer, and for a given high-Z material dissolved in the atmosphere by ablation, its local abundance determines its exact partial pressure in the gas. If the partial pressure of each high-Z material exceeds its saturated vapor pressure, which is a function of the ambient temperature in that layer, then this means that the high-Z material is in excess of a saturated state, and it must condense, releasing latent heat during the phase transition. Condensed matter descends towards the core if it is heavier than the ambient gas, losing gravitational energy on its way. This so-called ’rainout’ process leads to a compositional gradient. Since the temperature increases with deepening layers, each layer is able to contain more vapor of the same material (the saturated vapor pressure increases with the temperature). At the boundary of the solid silicate core and the overlying saturated atmosphere, if the latter cannot absorb all the (continually) ablated high-Z material, the rest of it adds to the solid core. This allows the solid core to grow even without direct impacts (Iaroslavitz and Podolak, 2007; Leconte et al., 2017; Bodenheimer et al., 2018; Brouwers et al., 2018; Brouwers and Ormel, 2020). The rainout process implies that high-Z materials increase inward according to Z. Given differences in their mean molecular weights, saturated vapor pressures and the temperature gradient in the atmosphere, one expects a stratified structure to emerge, just as it does in our growth and evolution code (Figures 6-8).
@@ 4.2 Chemistry During Accretion
Our outer solar system planetesimals are represented by the following simplified components: FeSiO (fayalite) represents the rocky mineral component; C (graphite) represents the organic (or idealized carbon-rich) refractory component; The icy components are HO and CO. While there are many possible mineralogies to choose from, the one that was chosen above has a specific intention. By adding hydrogen, we can chemically separate FeSiO into Fe, SiO and HO, while graphite will convert to CH, and CO to CH and HO, and therefore all of these chemical products (iron,silica,water,methane and envelope) are standard materials that are most frequently used in canonical models of gas giants, and have well tested equations of state. We maintain flexibility in the model, by varying the ratios among the reactants.
@@
For chemical equilibrium calculations we utilize the widely used STANJAN code (Reynolds, 1986), which provides the chemical equilibrium state of an ideal gas mixture. The list of elements that can be employed in STANJAN is quite extensive, however Fe is not one of the included elements in the code. Therefore, of our three reactions, only that of fayalite decomposition could not be accounted for with STANJAN. Fortunately, the fayalite decomposition experiments robustly cover a much wider range of T-P conditions compared to the other two reactions, spanning over 5 orders of magnitude in pressure (Massieon et al., 1993; Efimchenko et al., 2019). As a first step we validate the two other experimental results for CO methanation (Johnson and Nuth, 2015), and C methanation (Peña-Alvarez et al., 2021). We find that STANJAN reproduces the same results as the experiments, when we input the same T-P conditions into the code. We therefore gain confidence in its predictive capabilities, while we plan to use more elaborate codes in future studies.
In order to investigate methanation for the entire range of T-P conditions shown in Table 1, we consider the elements C, H and O. Of course, since Fe and Si are excluded from STANJAN, the chemical equilibrium resulting in these calculations gives an incomplete picture, and is only designed to obtain a partial view of the chemistry. Using these elements we can account for the five species that participate in Equations 2 and 3, namely: H, C, CO, HO and CH. We require the relative fraction of each species in a state of equilibrium at constant pressure and temperature, but also for completion we look at several other potential chemical products which might stably form at the appropriate conditions: CO (carbon dioxide), CH (ethane) and CHOH (methanol). Since we must neglect fayalite, we assume as a heuristic approach, that our input planetesimals consist of twice as much C (40% mass fraction) as HO (20% mass fraction), and equal mass fraction of HO and CO (20%). Our input mass fraction for H (20%) is selected arbitrarily in order to provide sufficient hydrogen for full chemical transformation of C, HO and CO. We note that our primary concern is not to explore various realizations of the input composition, but rather to investigate if the chemistry goes in the direction of methane or not. For this purpose, small changes in the initial composition do not qualitatively alter the outcomes. We demonstrate the effect of the temperature and pressure state on the chemistry in Table 2, where we show the relative mass fractions of the various chemical products, given a range of conditions.
| T-P state \ Species | H | C | CO | HO | CH | CO | CH | CHOH |
@@
We can divide the chemical processes that act on high-Z materials into three distinct stages in the evolution: initial ablation of planetesimals, their subsequent rainout and the long-term global evolution of the planet. For the first stage, we use our planetesimal ablation code. Thus, we consider the values listed in Table 1 to represent the relevant conditions. Cross referencing these values with Table 2, it is clear that pebbles and small planetesimals are dissolved in the atmosphere early-on and experience conditions that drive the chemistry towards methane and water. Large planetesimals on the other hand might break up due to ram-pressure at locations where the conditions are more inclined to drive the chemistry towards less CH and more CO (and even CO), while HO also diminishes.
The second stage is rainout of high-Z materials towards the core. This process is included in our growth code for Uranus, however without explicitly accounting for chemistry, and only for mono-sized large planetesimals. We can nevertheless use it to approximately estimate the possible range in T-P conditions in the region of the atmosphere that contains hydrogen, as marked by the arrows in Figures 6-8. Increase in both temperature and pressure during rainout may result in the chemistry moving either towards methane or carbon monoxide. It is hard to determine the outcome without including explicit feedbacks in the code.
@@ 4.3 Interior Composition After Methane Formation
As discussed above, the conditions are likely appropriate for chemistry to occur in the growing atmosphere of Uranus and Neptune. Demonstrating that chemical conversion is complete remains a future task for much more advanced models than currently exist in the literature. We note that even if the chemical transformation is incomplete, it still applies to a very large fraction of the accreted solids. Since modeling of partial chemical evolution is extremely difficult, we instead simplify the problem by assuming, without full justification, that the chemical reactions we invoke apply to all solids. The methane fraction resulting from the calculation below may thus be viewed as an upper limit.
@@
In Figure 9 all the models in Table 3 (abundant hydrogen) are plotted to show the scatter in possible interior composition, against the background of randomly generated models. Using the most recent moment of inertia factor estimates for Uranus (0.22) and Neptune (0.24) (Neuenschwander and Helled, 2022) each random model is depicted by a single pixel, whose colour corresponds to the central temperature. The mass fraction of methane in the planet is shown as a function of the water/rock mass ratio (here rock = SiO+Fe). The dashed vertical line denotes equal proportions of water and rock. The right (left) dotted vertical line indicates a water/rock mass ratio of 2 (0.5). Superimposed on the same plot are the calculated upper-bound methane mass fractions based on chemistry, when accounting for the range in our reactants ratios (red asterisks). The composition derived from our simple chemical model falls well within the allowed parameter space of internal structures determined by our random models, independently producing the same patterns for both planets. In these models with abundant methane we see that the proportions between rock and water in Uranus and Neptune can shift either towards rock or water, but since methane itself is an ice, they are strictly always icy, despite having been initially formed by refractory-dominated solids. This is better reflected in the ice/rock mass ratio in Figure C1 in Appendix C. We likewise confirm that iron in our random models is within the expected range in solar system planets (Figure C2 in Appendix C).

### v2

@@ 3.4 Models with Methane
We also consider a third ice, methane, as motivated by the following arguments. The CH4 abundance in the upper atmosphere of Uranus and Neptune is observed to be far greater than solar (Atreya et al., 2020). This should not be regarded as indicative of the bulk abundance of methane, since these measurements only probe the atmosphere at the 1 bar level. Nevertheless, early studies did in fact consider models where CH4 was mixed into the entire envelope with far greater abundances than solar (Podolak, 1976; Podolak and Reynolds, 1981), or with an icy layer that consisted primarily of H2O and CH4 (Hubbard and Macfarlane, 1980). This was motivated by the idea that nebular carbon should condense to form CH4 and accrete onto Uranus and Neptune in large amounts. At around the same time, however, the view was shifting away from condensation of solid CH4 as a result of kinetic inhibition (Lewis and Prinn, 1980). Instead, at sufficiently low temperatures, the expectation was that the large majority of carbon would condense directly to CO (Lewis and Prinn, 1980).
Later it was briefly suggested that, instead, CO might be chemically transformed in the atmosphere of the gas giants and convert to CH4 (Podolak and Reynolds, 1984) when reacting with hydrogen, altering their bulk composition. However, this hypothesis did not develop further since. Here we extend the idea that CH4 could be a key volatile in the outer planets. As already outlined in Section 2, while CO is almost negligible and therefore has a very small potential contribution for generating methane in typical contemporary planetesimals, refractory organics contribute to a very significant portion of their mass. These refractory organics may also chemically transform into copious amounts methane when they accrete in vast quantities onto the planet’s atmosphere, and are capable of altering its bulk composition.
We use the same setup as in the previous section, but now replace carbon monoxide with methane. When we incorporate CH4 as our second icy material, the results are very different. Randomly generated interior models of Uranus and Neptune with CH4 are shown in Figure 3. In these models the water/rock mass ratio now spans the range 0.1-10. As noted above, CO as an additional ice tends to replace SiO2. CH4, on the other hand, tends to replace water, so the rock/water mass ratio can be significantly larger.
@@
Finally, in Figure 5, we show the methane mass fraction as a function of the water/rock mass ratio. The methane mass fraction probed by the space of possible solutions is at least 10%. In models that also have a large rock/water mass fraction (left of the vertical dashed line) it is necessarily larger, upwards of 20%. Since the envelope mass fraction in these models also has large values (Figure 3), the combination of large envelope and even larger CH4 mass fractions means that the other materials - H2O, SiO2, and Fe, are collectively only a relatively small fraction of the planet’s mass. Thus, in models that have a large rock/water mass fraction, specifically, the methane fraction might even be larger than that of H2O, SiO2, and Fe combined. This is a difficulty because CH4 is certainly not that prevalent in the contemporary solar system, and is normally a very small fraction compared to water (Bockelée-Morvan and Biver, 2017). Thus, for methane to participate in solving the composition tension problem, a large methane fraction must be justified.
We note that since in our current model we only mix two adjacent materials, it might be possible that other mixing choices would result in less methane. For instance, triple mixing of water, methane and H+He in the outer atmosphere might lower the bulk methane fraction. On the other hand, it seems reasonable to expect that if our current water + methane region and methane + envelope region were to be mixed into one homogeneous region, the overall material masses would not change dramatically. It remains a task for future studies to quantitatively explore the triple mixing possibility, but the methane fraction will likely remain too large to reconcile with its contemporary abundance in the solar system, either way.
We suggest instead that organic-rich refractories are sufficiently abundant in outer solar system planetesimals to drive chemical reactions in the atmospheres of Uranus and Neptune, which, during the phase of planet growth, may produce the required copious amounts of methane. These processes are discussed next.
@@ 4 Growth of the Outer Planets and Accretion of Solids
We consider a growing planet in the outer solar system, which starts accreting mass onto an embryo-sized core (the incipient core mass in the range 0.05-0.5 M⊕ (Brouwers et al., 2018; Bodenheimer et al., 2018)). Classical models emphasized subsequent growth due to accretion of large planetesimals. These planetesimals could fall directly onto the core (Podolak et al., 1988) however they could also dissolve in the atmosphere due to thermally (Iaroslavitz and Podolak, 2007; Brouwers et al., 2018) or mechanically (Mordasini et al., 2015) driven ablation. As core mass accumulates, additional atmosphere is collected (Pollack et al., 1996). Energy release is added to the luminosity of the planet and thus internal temperatures increase during growth (Pollack et al., 1996).
@@
We can get an idea of the relevant density profile, as well as the temperature-pressure (T-P) profiles, using a newly updated code (Stevenson et al., 2022) implemented here for the first time for a growing Uranus. This code solves the equations of mass conservation, hydrostatic equilibrium, energy conservation, and energy transport assuming spherical symmetry. It includes the effects of mass and energy deposition by accreting solids and gases, and equations of state of water, SiO2 and iron. It also includes opacity due to dust grains, the ablation and breakup of (here only 100 km mono-sized) solid planetesimals in the gaseous envelope, treatment of compositional gradients and heating by the host star which drives mass loss. Using this code, we obtain snapshots of the density/temperature/pressure profiles of the atmosphere, at three different times after the start of the evolution: 1.35, 2.06 and 2.65 Myr (see Figure 6).
We then use these atmospheric density profiles as input to an ablation code (Podolak et al., 1988), which includes the detailed effects of frictional gas drag, heating by ambient thermal radiation, and the ensuing evaporation (sublimation) erosion derived from the energy input. Following the trajectories of accreted planetesimals of various sizes, we can determine the characteristic temperature-pressure conditions during ablation. For the parameters we selected, we find that small sub-km planetesimals are strongly slowed by atmospheric drag and ablate as a consequence. The larger planetesimals are correspondingly less affected by drag and are therefore able to maintain the speed necessary to break up deeper in the atmosphere, due to ram pressure. As already mentioned, these results depend on the assumed material properties. Since the code is unable to account for the intricacies of planetesimal intra-porosity or material mixing, we employ a similar approach to other studies (Iaroslavitz and Podolak, 2007; Brouwers et al., 2018; Brouwers and Ormel, 2020; Ormel et al., 2021) and simplify matters by considering non-porous planetesimals with a uniform density of 2000 kg m-3, and the material thermodynamic/cohesion properties of water, which might represent a mean value for a mixture of silicates, organics, water and other high-volatility ices. Given these assumptions, the temperature-pressure conditions during ablation are provided in Table 1. Note that ablation is a gradual process and therefore it occurs for a range of T-P conditions. To simplify matters, for sub-km planetesimals we listed a single T-P value that matches the moment during which half of the planetesimal mass has eroded. For larger planetesimals, the listed value matches the moment of ram-pressure breakup when most of the mass is instantantly delivered to the atmosphere.
@@ 4.1 Compositional Stratification During Accretion
Once high-Z materials are present in the atmosphere, they are expected to form a compositional gradient inside the planet (Helled and Stevenson, 2017; Lozovsky et al., 2017; Brouwers et al., 2018; Bodenheimer et al., 2018; Valletta and Helled, 2019; Vazan and Helled, 2020; Brouwers and Ormel, 2020; Ormel et al., 2021). Indeed there is some seismic and gravity-based evidence for a deep compositional gradient in Saturn (Fuller, 2014; Mankovich and Fuller, 2021) and Jupiter (Wahl et al., 2017; Debras and Chabrier, 2019; Howard et al., 2023).
Assuming that high-Z vapor in the atmosphere obeys the ideal gas law, an assumption which was found to be judicious in giants such as Uranus and Neptune (Iaroslavitz and Podolak, 2007), then compositional stratification can be explained as follows. At a given layer, and for a given high-Z material dissolved in the atmosphere by ablation, its local abundance determines its exact partial pressure in the gas. If the partial pressure of each high-Z material exceeds its saturated vapor pressure, which is a function of the ambient temperature in that layer, then this means that the high-Z material is in excess of a saturated state, and it must condense, releasing latent heat during the phase transition. Condensed matter descends towards the core if it is heavier than the ambient gas, losing gravitational energy on its way. This so-called ’rainout’ process leads to a compositional gradient. Since the temperature increases with deepening layers, each layer is able to contain more vapor of the same material (the saturated vapor pressure increases with the temperature). At the boundary of the solid silicate core and the overlying saturated atmosphere, if the latter cannot absorb all the (continually) ablated high-Z material, the rest of it adds to the solid core. This allows the solid core to grow even without direct impacts (Iaroslavitz and Podolak, 2007; Leconte et al., 2017; Bodenheimer et al., 2018; Brouwers et al., 2018; Brouwers and Ormel, 2020). The rainout process implies that high-Z materials increase inward according to Z. Given differences in their mean molecular weights, saturated vapor pressures and the temperature gradient in the atmosphere, one expects a stratified structure to emerge, just as it does in our growth and evolution code. The composition distribution for the fiducial materials iron, silica, water and H+He is shown in Figure 6.
| (a) 1.35 Myr |
| (b) 2.06 Myr |
| (c) 2.65 Myr |
@@ 4.2 Chemistry During Accretion
Our outer solar system planetesimals are represented by the following simplified components: Fe2SiO4 (fayalite) represents the rocky mineral component; C (graphite) represents the organic (or idealized carbon-rich) refractory component; The icy components are H2O and CO. While there are many possible mineralogies to choose from, the one that was chosen above has a specific intention. By adding hydrogen, we can chemically separate Fe2SiO4 into Fe, SiO2 and H2O, while graphite will convert to CH4, and CO to CH4 and H2O, and therefore all of these chemical products (iron,silica,water,methane and envelope) are standard materials that are most frequently used in canonical models of gas giants, and have well tested equations of state. We maintain flexibility in the model, by varying the ratios among the reactants.
@@
For chemical equilibrium calculations we utilize the widely used STANJAN code (Reynolds, 1986), which provides the chemical equilibrium state of an ideal gas mixture. The list of elements that can be employed in STANJAN is quite extensive, however Fe is not one of the included elements in the code. Therefore, of our three reactions, only that of fayalite decomposition could not be accounted for with STANJAN. Fortunately, the fayalite decomposition experiments robustly cover a much wider range of T-P conditions compared to the other two reactions, spanning over 5 orders of magnitude in pressure (Massieon et al., 1993; Efimchenko et al., 2019). As a first step we validate the two other experimental results for CO methanation (Gao et al., 2012), and C methanation (Peña-Alvarez et al., 2021). We find that STANJAN reproduces the same results as the experiments, when we input the same T-P conditions into the code. We therefore gain confidence in its predictive capabilities, while we plan to use more elaborate codes in future studies.
In order to investigate methanation for the entire range of T-P conditions shown in Table 1, we consider the elements C, H and O. Of course, since Fe and Si are excluded from STANJAN, the chemical equilibrium resulting in these calculations gives an incomplete picture, and is only designed to obtain a partial view of the chemistry. Using these elements we can account for the five species that participate in Equations 2 and 3, namely: H2, C, CO, H2O and CH4. We require the relative fraction of each species in a state of equilibrium at constant pressure and temperature, but also for completion we look at several other potential chemical products which might stably form at the appropriate conditions: CO2 (carbon dioxide), C2H6 (ethane) and CH3OH (methanol). Since we must neglect fayalite, we assume as a heuristic approach, that our input planetesimals consist of twice as much C (40% mass fraction) as H2O (20% mass fraction), and equal mass fraction of H2O and CO (20%). Our input mass fraction for H2 (20%) is selected arbitrarily in order to provide sufficient hydrogen for full chemical transformation of C, H2O and CO. We note that our primary concern is not to explore various realizations of the input composition, but rather to investigate if the chemistry goes in the direction of methane or not. For this purpose, small changes in the initial composition do not qualitatively alter the outcomes. We demonstrate the effect of the temperature and pressure state on the chemistry in Table 2, where we show the relative mass fractions of the various chemical products, given a range of conditions.
| T-P state \ Species | H2 | C | CO | H2O | CH4 | CO2 | C2H6 | CH3OH |
@@
We can divide the chemical processes that act on high-Z materials into three distinct stages in the evolution: initial ablation of planetesimals, their subsequent rainout and the long-term global evolution of the planet. For the first stage, we use our planetesimal ablation code. Thus, we consider the values listed in Table 1 to represent the relevant conditions. Cross referencing these values with Table 2, it is clear that pebbles and small planetesimals are dissolved in the atmosphere early-on and experience conditions that drive the chemistry towards methane and water. Large planetesimals on the other hand might break up due to ram-pressure at locations where the conditions are more inclined to drive the chemistry towards less CH4 and more CO (and even CO2), while H2O also diminishes.
The second stage is rainout of high-Z materials towards the core. This process is included in our growth code for Uranus, however without explicitly accounting for chemistry, and only for mono-sized large planetesimals. We can nevertheless use it to approximately estimate the possible range in T-P conditions in the region of the atmosphere that contains hydrogen, as marked by the arrows in the various panel of Figure 6. Increase in both temperature and pressure during rainout may result in the chemistry moving either towards methane or carbon monoxide. It is hard to determine the outcome without including explicit feedbacks in the code.
@@ 4.3 Interior Composition After Methane Formation
As discussed above, the conditions are likely appropriate for chemistry to occur in the growing atmosphere of Uranus and Neptune. Demonstrating that chemical conversion is complete remains a future task for much more advanced models than currently exist in the literature. We note that even if the chemical transformation is incomplete, it still applies to a very large fraction of the accreted solids. Since modeling of partial chemical evolution is extremely difficult, we instead simplify the problem by assuming, without full justification, that the chemical reactions we invoke apply to all solids. The methane fraction resulting from the calculation below may thus be viewed as an upper limit only, since (a) we assumed complete chemical conversion; and (b) organic refractories are dominated by carbon, but are not 100% carbon like we envisioned in our simple chemical scheme.
@@
In Figure 7 all the models in Table 3 (abundant hydrogen) are plotted to show the scatter in possible interior composition, against the background of randomly generated models. Using the most recent moment of inertia factor estimates for Uranus (0.22) and Neptune (0.24) (Neuenschwander and Helled, 2022) each random model is depicted by a single pixel, whose colour corresponds to the central temperature. The mass fraction of methane in the planet is shown as a function of the water/rock mass ratio (here rock = SiO2+Fe). The dashed vertical line denotes equal proportions of water and rock. The right (left) dotted vertical line indicates a water/rock mass ratio of 2 (0.5). Superimposed on the same plot are the calculated upper-bound methane mass fractions based on chemistry, when accounting for the range in our reactants ratios (red asterisks). The composition derived from our simple chemical model falls well within the allowed parameter space of internal structures determined by our random models, independently producing the same patterns for both planets. In these models with abundant methane we see that the proportions between rock and water in Uranus and Neptune can shift either towards rock or water, but since methane itself is an ice, they are strictly always icy, despite having been initially formed by refractory-dominated solids. This is better reflected in the ice/rock mass ratio in Figure C1 in Appendix C. We likewise confirm that iron in our random models is within the expected range in solar system planets (Figure C2 in Appendix C).